x + 1 2. Evaluate the following integral. This involves substitution and integration of In. Si
Answer #1

3. x + 1 da O It å asked to evaluate x 5452 . Let u= x 3 + 50 du ax (25+5x] da - 52 +5 = 5(**+1) - du = (x + 1;)dx when x = 2حل x² + I do I lo 43 D 25+5% Note S4 I dx dx = Inlal + c e-constant of integration n-1 dan NX dx • In(a) – Inb) = in